You are given the data below for 2008 for the imaginary country of Amagre, whose currency is the G.


Consumption 350 billion G

Transfer payments 100 billion G

Investment 100 billion G

Government purchases 200 billion G

Exports 50 billion G

Imports 150 billion G

Bond purchases 200 billion G

Earnings on foreign investments 75 billion G

Foreign earnings on Amagre investment 25 billion G


Compute net foreign investment.

Compute net exports.

Compute GDP.

Compute GNP.


1
Expert's answer
2021-03-04T10:36:47-0500

Net foreign investment = Net exports


Net export = Export - Import

 Net exports = 50 - 150 = -100 billion G.


Net foreign investment = 100 billion G.


GDP is equal to the sum of consumption, government spending, investment and net exports.

GDP = C + I + G + NX

GDP = 350 + 100 + 200 - 100 = 550 billion G.


GNP = C + I + G + NX + NY

GNP = GDP + NY

GNP = 550 + (25 - 75) = 500 billion G.